OpenStudy (anonymous):

the gravitational force between two objects varies inversely as the square of the distance between the objects. if a force of 30 units results from two objects that are 12 units apart, how much force results from two objects that are 15 units apart?

OpenStudy (anonymous):

30 = x/12^2. thus x = 30*144. then ? = x/15. ? = 30*144/15=2*144 = 288 force units.

OpenStudy (anonymous):

288/9

OpenStudy (anonymous):

thank you guys